OpenStudy (anonymous):

Find the exponential Inequality 5^(2x + 2 )< 625

OpenStudy (anonymous):

to solve this, you first take the natural log of both sides \[\ln(5^{2*x+2})<\ln(625)\] Using laws of logarithms, \[(2x+2)*\ln(5)<\ln(625)\] solving for x \[x<({\ln(625)/\ln(5)}-2)/2\]

OpenStudy (anonymous):

\[5^{2x+2}\leq 5^4\iff 2x+2\leq 4\] no need for logs

OpenStudy (anonymous):

\[2x+2\leq 4\] \[2x\leq 2\] \[x\leq 1\]

OpenStudy (anonymous):

the only thing you need to know for this problem is that \(5^4=625\)